Prove using mathematical induction: (4) Prove that for all n E N, 3(7 – 4”).
Answer #1

to show that, 31 (7-4). step 1: to show that result holds for n=1. for n=1 70-40 = 7-4 = 3 AS 313 » Result holds for nal.